OpenStudy (anonymous):

The equation of a line is 2x-2y=6. What is the slope of the line?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: solve for y

OpenStudy (anonymous):

y=0?

jimthompson5910 (jim_thompson5910):

example: 5x+10y = 20 5x+10y-5x = 20-5x 10y = -5x+20 10y/10 = (-5x+20)/10 y = -1/2x + 2 The slope here is -1/2. This is not the answer as it's just an example

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

how are you getting that

OpenStudy (anonymous):

Never mind. That doesn't make since.

jimthompson5910 (jim_thompson5910):